Fermi National Laboratory


Accelerator Update for July 25, 2005 - July 27, 2005

Monday July 25
The day shift began with Operations monitoring store 4296, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and Meson MCenter experiments taking beam.

Operations began at 11:30 AM the Booster and Main Injector (MI) turn off in preparation for accesses and RF work. Pbar will maintain its stack, but the Recycler, MiniBooNE, NuMI, and Switchyard (SY) beam will all be off.

Operations began preparing NuMI for access at 11:50 AM.

Operators had the MI and Booster tunnels interlocked by 6:13 PM.

Operations established beam to Booster at 8:30 PM.

Operations established beam to MiniBooNE at 9:16 PM.

Operations established beam to Pbar and NuMI at 9:43 PM.

At 9:44 PM, a MI RF station (MIRF5) was tripping off. Operators reduced the number of turns to keep the station on. BRF3 was also tripping off.

Operations established beam to MCenter at 10 PM.

Tuesday July 26
The midnight shift began with Operations monitoring store 4296, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:27 AM, a Linac RF station (KRF2) tripped off and wouldn’t reset. Operations called in an expert. He got the station running and Pbar resumed stacking at 1:37 AM.

At 1:44 AM, BRF3 and BRF12 tripped off. They both reset okay.

At 1:47 AM, MIRF5 tripped off for the first of more times than I can count over the next six hours. It reset each time.

Operations put the Recycler into shot setup at 1:47 AM to stash antiprotons.

At 2:16 AM, NuMI tripped off due to a helium flow alarm. An operator called an expert who said that the alarm limit could be raised.

The Recycler stash ended and Pbar resumed stacking 2:55 AM.

MCenter resumed taking beam at 3:07 AM.

At 3:22 AM, BRF3 tripped off for the first of seven times in four hours. It reset each time.

At 4:08 AM, BRF12 & 15 tripped off. They both reset okay.

At 8:21 AM, MIRF5 tripped off several times in two minutes. Operations contacted an expert.

Operations terminated store 4296 at 12:08 PM. Operations prepared CDF and D0 for collision hall accesses.

Operations established beam to MTest at 12:39 PM.

The collision hall accesses were completed and operators began the TeV turn on.

Operations put the TeV through a dry squeeze at 3:42 PM.

Operations put the TeV into shot setup at 4:32 PM.

At 7:14 PM, after the protons and antiprotons had been loaded into the TeV, the TeV suffered a quench. Operations called in experts to investigate.

Pbar resumed stacking at 7:48 PM.

TeV experts reported at 9:55 PM that a separator may have sparked and caused the quench.

Operations put the TeV through a dry squeeze at 11:43 PM.

Wednesday July 27
The midnight shift began with the TeV separators in conditioning mode, with Operations monitoring a small stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations loaded the TeV with a 4x0 study-store at 1:36 AM. The study lasted until 4:06 AM.

Operations put the TeV into shot setup at 4:25 AM.

At 6:36 AM, a communication problem with the TeV sector A0 cryo system cold compressors put shot setup on hold. An expert came in and replaced a crate power supply at 7:47 AM

Recycler experts used their small stash for studies beginning around 7 AM.

At 7:09 AM, Operations reported losing the 4th shot of antiprotons. Operators investigated.

Operations reported that Pbar successfully transferred a small shot of antiprotons (5mA) into the TeV at 7:21 AM.

Operations established store 4301 at 7:58 AM with an initial luminosity of 56.36E30.

The Plan for Wednesday
The plans for today are to allow some machine studies and to conduct a Recycler stash. The store will remain in until the midnight shift on Thursday.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/27/2005   email Fermilab